OpenStudy (anonymous):

HELP ME PLEASE. Commonwealth Edison Company supplies electricity to residential customers for a monthly customer charge of $7.58 plus 8.295 cents per kilowatt-hour for up to 400 kilowatt-hours. (Hints: Pay attention to the units.) a.Write an equation that relates the monthly charge C, in dollars, to the number x of kilowatt-hours used in a month, 0≤x≤400. b.What is the monthly charge for using 150 kilowatt-hours?

OpenStudy (texaschic101):

a.) 7.58 + 8.295(k < 400) = C (I am not sure about this one) b.) 7.58 + 8.295(150) = total cost can you finish this one now ?

OpenStudy (anonymous):

oh ok..thanks for your attempt .@texaschic101

OpenStudy (texaschic101):

can you finish the 2nd one or do you need help ?

OpenStudy (anonymous):

i would only need to multiply 8.295 * 150 and 7.58

OpenStudy (anonymous):

add all through and dat would be my answer

OpenStudy (texaschic101):

hold on....the 8.295 is in cents... it might be .08295(150) = 12.44 + 7.58 = 20.02

OpenStudy (texaschic101):

@amistre64 ...can you check this please

OpenStudy (anonymous):

oh ok..i think you right now.was jux about to ask you if i dont need to convert the cents into dollars

OpenStudy (anonymous):

@texaschic101 should i write out the equation as 7.58 + 0.08295x

OpenStudy (texaschic101):

well...I know if you just write it out like this : 7.58 + 8.295(150) will give you the wrong answer because it looks like the 8.295 is dollars and not cents. .08295......is definitely cents.....but you might want to get a second opinion.

OpenStudy (anonymous):

OH OK..THANKS VERY MUCH..MUCH APPRECIATED